OpenStudy (anonymous):

Need help proving that the second derivative of y=a/2(e^(x/a)+e^(-x/a)) is y/a^2 through differentiation. Thanks

OpenStudy (accessdenied):

Well if you take the derivative of an exponential function, what happens? Taking derivative of e^(kx), you just get e^(kx) * k, right? So what we could do is differentiate the current function twice using exponents rule. Then factor some constants out to get the statement for y, along with the constant multiple 1/a^2 that we factor out.

OpenStudy (anonymous):

can you please help me with the first differentiation and i will then do the second

OpenStudy (accessdenied):

Sure. To make it easier we might distribute a/2 to both terms first... \( y = \dfrac{a}{2} \left( e^{x/a} + e^{-x/a} \right) = \dfrac{a}{2} e^{x/a} + \dfrac{a}{2} e^{-x/a} \)

OpenStudy (accessdenied):

Now when we take the derivative of both sides to find first derivative, we simply take the derivative term-by-term: Derivative of \( \dfrac{a}{2} e^{x/a} \). Note that the exponent \(x/a\) is just x multiplied by a constant, \(1/a\). So we use derivative of exponent \(\dfrac{d}{dx} e^{kx} = k e^{kx} \) \( \dfrac{d}{dx} \left( \dfrac{a}{2} e^{x/a} \right) = \dfrac{a}{2} e^{x/a} \times \dfrac{1}{a} \) Is any of this unclear so far?

OpenStudy (anonymous):

no i understand this thank you

OpenStudy (accessdenied):

Alright. Because every step after this is virtually identical. The derivative of the second term is nearly the same as the first, only our constant is -1/a instead of 1/a. \( \dfrac{d}{dx} \left( \dfrac{a}{2} e^{-x/a} \right) = \dfrac{a}{2} e^{-x/a} \times \left(-\dfrac{1}{a}\right) \)

OpenStudy (anonymous):

yes as it is just the difference of the negative

OpenStudy (accessdenied):

Giving us the first derivative: \(y' = \dfrac{a}{2} e^{x/a} \times \dfrac{1}{a} + \dfrac{a}{2} e^{-x/a} \times \dfrac{-1}{a} \) You could simplify but what we end up doing is factoring out the 1/a^2 again to show that it is y/a^2. :)

OpenStudy (accessdenied):

Factoring out the 1/a^2 after we take the second derivative, that is *

OpenStudy (anonymous):

ok

OpenStudy (accessdenied):

Were you able to figure it out from there?

OpenStudy (anonymous):

i beleive so

OpenStudy (accessdenied):

Alright. Essentially if you obtained the result: \(y''= \dfrac{a}{2} e^{x/a} \times \dfrac{1}{a} \dfrac{1}{a} + \dfrac{a}{2} e^{-x/a} \times \dfrac{-1}{a} \dfrac{-1}{a} \) \(y''= \dfrac{1}{a^2} \underbrace{\left( \color{blue}{\dfrac{a}{2} e^{x/a} + \dfrac{a}{2} e^{-x/a}}\right) }_{y} \) \( y''= y / a^2 \) You would be golden. :)

OpenStudy (anonymous):

thank you so much i almost had it

OpenStudy (accessdenied):

Glad to help! :)